the point charges Q and -2Q are placed at some distance apart. if the electric field at the location of Q is E. then,the electric field at tge location of -2Q will

Grade:12th pass

1 Answers

Eshan
5 years ago
Dear student,

Electric field at location of Q is due to -2Q. Let the distance between them bed.

Therefore, electric field at Q= E = \dfrac{k(2Q)}{d^2}
Electric field at -2Q is due to Q only and equat to\dfrac{kQ}{d^2}=\dfrac{E}{2}